The directions are as follows: If c is the measure of the hypotenus, find each missing measure. Round to the nearest 10th if necessary.
a=?, b=15, c=31 I understand how to use the Pythagorean theorem to work problems where c is missing, but here I am confused. I don't even know where to begin.

Answer by edjones(8007)

a^2+15^2=31^2
a^2=31^2-15^2
a=sqrt(31^2-15^2)
a=27.1
.
Ed